OpenStudy (anonymous):

5x-7(2-4x) And can you explain how you got the answer

OpenStudy (anonymous):

FOIL / Distribute Leave 5x alone for now. -7(2-4x) = -7*2 and -7*-4x = -14+28x Add in 5x -> 5x - 14 + 28x

OpenStudy (anonymous):

But what is the rule when you do it for any problem like that

OpenStudy (anonymous):

You expand the equation in any way possible until it cannot be expanded. -7(2-4x) can be expanded into -14+28x through distribution. After that, nothing else can be expanded.

OpenStudy (anonymous):

Ohhh Thank you